The volume of O2 liberate from 0.96 of H2O2 at STP solve the questions

Grade:11

1 Answers

Arun
6 years ago
no.of moles of  H2O2 is= 0.96/68=0.014
(68=molecular mass of hydrogen peroxide)
H2O2===>H2 + O2
hence volume of O2 evolved =22.4L*0.014=0.3136L=313.6 mL (22.4L=volume of 1 mole of ideal gas)
 
Regards
Arun (askIITians forum expert)